x+9 = x-10
---- ----
5 11





X+9 = X-10

Answer by checkley77(12844)


x+9 = x-10
---- ----
5 11
11(X+9)=5(X-10)
11X+99=5X-50
11X-5X=-50-99
6X=-149
X=-149/6 ANS.
PROOF:
[(-149/6)+9]/5=[(-149/6)-10]/11
11[(-149/6)+9]=5[(-149/6)-10]
-1,639/6+99=-745/6-50
-1,639/6+745/6=-50-99
-894/6=-149
-149=-149